OpenStudy (emeri_renteria):

what is the sum of 5/6 + 2/3, plzzzz help!!!!!!

OpenStudy (emeri_renteria):

plz somone help

OpenStudy (tkhunny):

Have you consdered a common denominator?

OpenStudy (mhchen):

(2/3) = (4/6) So you can rewrite like (5/6) + (4/6)

OpenStudy (emeri_renteria):

ohh k

OpenStudy (emeri_renteria):

but i need it as an mixed number so ima have to multiply and add

OpenStudy (emeri_renteria):

or i just divide?

Awolflover1 (awolflover1):

yes \[9\div6=? \]

OpenStudy (emeri_renteria):

1.5

Awolflover1 (awolflover1):

correct (: now you know to to convert Decimal to Fraction?

OpenStudy (tkhunny):

You can't be serious. You're using decimals to get to mixed numbers? You already had fractions. You need fractions. Don't let them go. \(\dfrac{5}{6} + \dfrac{2}{3} = \dfrac{5}{6} + \dfrac{4}{6} = \dfrac{9}{6} = \dfrac{3}{2} = \dfrac{2}{2} + \dfrac{1}{2} =\) 1½ Why would you use decimals to get there?

OpenStudy (mathmale):

Because both fractions in "(5/6) + (4/6)" have the same denominator, the "lowest common denominator," you can add them together. Add together their numerators, but keep the LCD (6). Your result is an "improper fraction," because your numerator is larger than your denom. If you want a mixed number, divide the denom. (6) into the numerator and write the result as \[1~\frac{ ? }{ 6 }\]

OpenStudy (mathmale):

If you haven't already finished this problem solving, do it now, please. I'm closing this post because it's been 13 days since we last heard from you.